OpenStudy (anonymous):

radical 5 times radical 8

geerky42 (geerky42):

Like what is shown in http://openstudy.com/study#/updates/5441e53be4b06e70924f7eca Since \(\sqrt{m}\sqrt{n} = \sqrt{m\cdot n}\), we have \(\sqrt5\sqrt8=\sqrt{5\cdot 8}=\sqrt{40}\) \(\sqrt{40} = \sqrt{4\cdot10}=\sqrt4\sqrt{10}=\boxed{2\sqrt{10}}\)

geerky42 (geerky42):

Clearly not 40.

OpenStudy (anonymous):

Yes i forgot to factor out